OpenStudy (anonymous):

PLEASE PLEASE PLEASE PLEASE PLEASE HELP a grocer wants to make a 10-pound mixture of peanuts and cashews that he can sell for $4.75 per pound., how many pounds of each should he use? let p=pounds of peanuts and c=pounds of cashews. write a system of equations that can be used to solve the problem

OpenStudy (anonymous):

@Squirrels

OpenStudy (squirrels):

$w@qq3r

OpenStudy (anonymous):

i can't access that link :(

OpenStudy (squirrels):

Let p be the pounds of peanuts Let c be the pounds of cashews Write an equation for the total weight. p + c = 10 Write an equation for the mixture. Remember the final mixture is $4.75 for each of 10 lbs so multiply 4.75 x 10: 4.00p + 6.50c = 4.75(10) Let's simplify the second equation: 4p + 6.5c = 47.5 Multiply both sides by 2 to get rid of the decimals: 8p + 13c = 95 Then solve the first equation for one variable (say p): p = 10 - c Substitute: 8(10-c) + 13c = 95 Distribute the 8 through the parentheses: 80 - 8c + 13c = 95 80 + 5c = 95 Subtract 80 from both sides: 5c = 15 Divide both sides by 5: c = 15/5 c = 3 Then substitute back to get p: p = 10 - c p = 10 - 3 p = 7 Answer: 7 lbs of peanuts 3 lbs of cashews

OpenStudy (squirrels):

$w@qq3r

OpenStudy (anonymous):

dang. thanks! i <3 you forever and ever

OpenStudy (squirrels):

<3